### v1

@@ 5 Conclusions
We have analyzed the influence of the selection procedure of most energetic particles itself in appearing of alignment phenomenon in the framework of HYDJET++ model. Event-by-event the transverse momentum conversation has been taken into account in the form of missing transverse momentum since the majority of particles are independently generated in the statistical model approach. It has been shown that the high degree of alignment appears at the reasonable values (up to $\sim 3$ GeV) of transverse momentum disbalance of selected most energetic particles. Applying this model we have the natural threshold on collision energy in appearing of alignment phenomenon since the conception of quark gluon plasma is valid after some minimal colliding energy. However one should note that the degree of alignment $P_{N}$ in our modeling is not enough large to match the central values of experimental measurements for $N\geq 4$ living some room for other explanations and further investigations.
@@ Acknowledgements.
It is pleasure to thank A.S. Chernyshov, A.I. Demianov, S.N. Nedelko for fruitful discussions. This work is supported by the Russian Science Foundation, grant 24-22-00011.

### v3

@@ V Discussion and Conclusions
Presumably, the alignment can manifest itself in both high energy collisions of relatively light nuclei (in cosmic rays experiments) and heavy ions (in collider experiments) together with collective effects, such as quark-gluon plasma formation and the local conservation of the transverse momentum of detected particles and their clusters. We have analyzed the influence of the selection procedure of most energetic particles itself in appearing of alignment phenomenon in the framework of HYDJET++ model. Event-by-event the transverse momentum conservation has been taken into account in the form of missing transverse momentum since the majority of particles are independently generated in the statistical model approach. The results on the figures 13 show that the high degree of alignment appears at the reasonable values (up to $\sim 3$ GeV) of transverse momentum disbalance of selected most energetic particles. In addition, the HYDJET++ model assumes the statistical particle production due to the hadronization of the quark-gluon plasma that automatically implies a collision energy threshold. This threshold is also characteristic of alignment observation.
Figure 4: Comparison of the simulation results of the alignment degree $P_{N}$ for three, four, five particles with the data of the Pamir collaboration. (a) – linear scale, (b) – logarithmic scale. “HJ++ with $\Delta$” – the modeling with taking account the local transverse momentum conservation with disbalance $\Delta$ (10); “HJ++” – the modeling without ${\bf p}_{T}$ conservation in every event. The values of the disbalance $\Delta$, which correspond to the our “best” alignment degree, are in the range of $0\div 1$ GeV.
A comparison of the simulation results with the Pamir collaboration data is shown in figure 4. It should be noted that since the Pamir data were obtained from experiments with cosmic rays, while our results are related to heavy ion collisions, rather an qualitative comparison can be made. Figure 4 shows that the alignment degree of three particles ($N=3$) can be fully described within the HYDJET++ model taking into account the local transverse momentum conservation with the disbalance $\Delta$, whereas for four ($N=4$) and five ($N=5$) particles we do not reproduce the experimental data [11] even taking into account measurement error:
$P_{3}^{\rm exp}=0.83\pm 0.27;~{}P_{4}^{\rm exp}=0.67\pm 0.33;~{}P_{5}^{\rm exp}=0.33\pm 0.23,$ (12)
while our “best” values are
$P_{3}=1;~{}P_{4}\approx 0.07;~{}P_{5}\approx 0.04,$ (13)
which achieved at disbalance in the interval
$\Delta=0\div 1~{}\text{GeV},$ (14)
corresponding to the transverse momentum conservation with high enough accuracy. Nevertheless the alignment degree $P_{4}$ and $P_{5}$ are in the limits of the two standard deviations from the experimentally measured values as a minimum, in addition, the values $P_{4}$ and $P_{5}$ increase by a factor of about 2 and 5, respectively, compared to result (9) due to ${\bf p}_{T}$ conservation in every event.
Thus our modeling demonstrates clearly that the conservation law of the transverse momentum via minimization of the disbalance $\Delta$ of the total transverse momentum of highest-energy particles allows one to select automatically more line up configurations. The application of this procedure to the HYDJET++ model can be partially justified by the fact that the majority of soft particles are generated independently as it was mention above and the transverse momentum conserves only in average, but in any rapidity interval. However one should note that the degree of alignment $P_{N}$ in our modeling is not enough large to match the central values of experimental measurements for $N\geq 4$ living some room for other explanations and further investigations.
@@ Acknowledgements.
It is pleasure to thank A.S. Chernyshov, A.I. Demianov for fruitful discussions and S.N. Nedelko for the our communication . This work is supported by the Russian Science Foundation, grant 24-22-00011.